OpenStudy (anonymous):

use the quadratic formula to solve the equation x^2+3=-4x

OpenStudy (anonymous):

bring over the -4x to the LHS. then your equation is of the form ax^2 + bx + c = 0 then use the quadratic formula.

OpenStudy (anonymous):

so what is the answer, i'm not good at doing this ?!

OpenStudy (anonymous):

\[x^2+4x+3=0\] \[(x+1)(x+3)=0\] \[x=-1\] or \[x=-3\]

OpenStudy (gw2011):

The quadratic formula is: x=[-b+/-square root (b^2-4ac)]/2a x^2+4x+3 a=1 because the coefficient of x^2 is 1 b=4 c=3 x=[(-4)+/-square root( 4^2-4*1*3)]/2(1) x=[(-4)+/-square root(16-12)]/2 x=(-4)+/-square (4)]/2 x=(-4+/-2)/2 x=(-4+2)/2=-2/2=-1 x=(-4-2)/2=-6/2=-3 Therefore, x=-1 and x=-3 You will notice that this problem can be solved without using the quadratic formula as satellite73 has shown. However, since your problem stated that you needed to solve this problem by using the quadratic formula, the above is the way of using the quadratic formula.